# Machine Learning Fleet Efficiency: Analyzing and Optimizing Large-Scale Google TPU Systems with ML Productivity Goodput

### v1

@@ Abstract.
Recent years have seen the emergence of machine learning (ML) workloads deployed in warehouse-scale computing (WSC) settings, also known as ML fleets. As the computational demands placed on ML fleets have increased due to the rise of large models and growing demand for ML applications, it has become increasingly critical to measure and improve the efficiency of such systems. However, there is not yet an established methodology to characterize ML fleet performance and identify potential performance optimizations accordingly. This paper presents a large-scale analysis of an ML fleet based on Google’s TPUs, introducing a framework to capture fleet-wide efficiency, systematically evaluate performance characteristics, and identify optimization strategies for the fleet. We begin by defining an ML fleet, outlining its components, and analyzing an example Google ML fleet in production comprising thousands of accelerators running diverse workloads. Our study reveals several critical insights: first, ML fleets extend beyond the hardware layer, with model, data, framework, compiler, and scheduling layers significantly impacting performance; second, the heterogeneous nature of ML fleets poses challenges in characterizing individual workload performance; and third, traditional utilization-based metrics prove insufficient for ML fleet characterization. To address these challenges, we present the “ML Productivity Goodput” (MPG) metric to measure ML fleet efficiency. We show how to leverage this metric to characterize the fleet across the ML system stack. We also present methods to identify and optimize performance bottlenecks using MPG, providing strategies for managing warehouse-scale ML systems in general. Lastly, we demonstrate quantitative evaluations from applying these methods to a real ML fleet for internal-facing Google TPU workloads, where we observed tangible improvements.
@@ 1. Introduction
This paper presents a playbook for instrumenting and optimizing large-scale ML fleets, using Google’s production TPU infrastructure for internal workloads as our case study. In contrast to prior work, where previous Google TPU papers have predominantly focused on hardware architecture design and chip-level performance evaluations (Jouppi et al., 2023, 2021, 2017), our work is a significant departure. We provide the first in-depth perspective on the software stack that enables these domain-specific accelerators to operate efficiently at scale.
[…]
Hardware Heterogeneity: Large-scale ML fleets typically comprise a mix of general-purpose processors (e.g., CPUs) and specialized accelerators (e.g., GPUs and TPUs). There is a strong interdependence between these diverse hardware components, as they often work in tandem to execute complex ML workloads. For instance, CPUs might handle data preprocessing and orchestration, while GPUs or TPUs perform the intensive matrix computations required for model training or inference. Each hardware type has its own performance characteristics, strengths, and weaknesses, and their effective coordination is crucial for overall system performance. For example, in our case at Google, within a single decade since the introduction of TPUs (Jouppi et al., 2017), ML fleets have grown to include multiple TPU variations, each with distinct features, configurations, and optimization strategies, all of which must be coordinated with the capabilities of the CPUs and other accelerators in the system. This hardware diversity and interdependence make it challenging to optimize fleet performance holistically.
Workload Heterogeneity: ML fleets must accommodate a variety of workloads and tasks, which evolve at a much faster pace than traditional WSC applications. This rapid evolution and diversity of workloads is driven by shifts in user demands for ML applications and changes in model architectures. New model types, such as large language models (Vaswani et al., 2017) or multi-modal systems (Driess et al., 2023; Yu et al., 2022), can dramatically alter the computational demands on the fleet.
ML fleets also consist of workloads from each stage of the ML lifecycle, encompassing training and inference phases. Each phase has its own unique computational characteristics and requirements. For example, training workloads demand high computational power and parallel processing capabilities (Kaplan et al., 2020), while real-time inference workloads prioritize low latency and high throughput (Pope et al., 2023).
The dynamic nature of ML workloads, coupled with the varying requirements across training and inference, demonstrates how the ML fleet must constantly adapt to new computational demands, balancing resources between different phases and workloads.
Hardware/Software Co-design: New ML workloads drive the development of new accelerators, while advances in hardware capabilities can enable novel ML applications. This tight coupling creates a moving target for optimization, as both the workloads and the hardware are in a state of constant evolution. The fast-paced ML landscape means that ML fleets must balance the need for specialization with the need for fungibility. For example, Figure 1 illustrates the explosive growth of DSAs in Google’s production fleet over a period of five years, driven by the rise of ML-based workloads.
Domain-specific hardware must be highly specialized to ensure maximal efficiency, but also flexible enough to quickly accommodate new workloads. This challenge is further compounded by the need for runtime systems, compilers, and schedulers to effectively bridge the gap between these evolving hardware capabilities and diverse workload requirements. The software stack must adapt to map workloads onto the most suitable hardware, optimize code for accelerators, and manage resources across the heterogeneous fleet.
The fundamental challenge is that all of these factors combine to create a complex, ever-changing landscape that makes it difficult to maintain optimal fleet performance. Traditional optimization strategies, which often rely on stable hardware and workload characteristics, are inadequate for addressing these dynamic interactions. Architecture-centric metrics, such as TOPs/Watt or peak FLOPS, while valuable for specific hardware evaluations, fall short in capturing the overall complexity and efficiency of an ML fleet.
[…]
2 Our second contribution addresses this gap. To view the fleet holistically and capture its performance across multiple layers of the ML system stack, we introduce the “ML Productivity Goodput” (MPG) metric. Unlike traditional metrics, MPG encompasses scheduling efficiency, runtime efficiency, and compiler / program efficiency. MPG is analogous to the “iron law” of computer performance (Eeckhout, 2010), but adapted for ML fleets as shown in Figure 2.
The MPG metric enables us to identify areas for improvement across the entire machine learning fleet by analyzing performance across different segments, such as accelerator type, model architecture, and workload phase (e.g., training vs. serving). By breaking down the metric into its components—Scheduling Goodput (SG), Runtime Goodput (RG), and Program Goodput (PG)—and examining it along the axes of fleet characteristics, we can precisely pinpoint potential optimization opportunities.
Figure 2. ML Productivity Goodput (MPG) and its components.
[…]
3 Finally, we provide a procedure for leveraging the MPG metric and demonstrate strategies that we have developed and deployed in the real world for optimizing ML fleet performance with the help of this metric. We show how MPG helps us identify compiler optimizations to generate more efficient code for specific hardware configurations, and improve scheduling algorithms to better utilize heterogeneous resources within the fleet, along with a few other strategies. Even after optimizations are implemented in the production fleet operations, the MPG metric can be used to validate and track fleet-wide improvements.
@@
In summary, we make the following contributions to advance the understanding and optimization of large-scale ML fleets:
  1. (1)
    An end-to-end description and dissection of the software and hardware stack required for a production ML fleet. We discuss the design choices that support large-scale ML workloads in a fleet consisting of domain-specific TPUs.
  2. (2)
    An examination of workload and hardware heterogeneity in a real ML fleet, revealing the challenges presented by evolving fleet demands and accelerator architectures.
  3. (3)
    Introduction of “ML Productivity Goodput” (MPG), a composable metric that tracks the performance of an ML fleet based on scheduling, runtime and compiler efficiency.
  4. (4)
    Strategies for optimizing fleet performance, through case studies and real-world applications of the MPG metric to identify and address bottlenecks across the system stack.
@@ 2.1. Workloads
WSCs (Barroso and Hölzle, 2009) traditionally focus on Internet-scale services, which serve web content, process user queries, and handle data storage and retrieval for billions of users. WSCs must manage diverse, often latency-sensitive tasks such as search indexing (Brin and Page, 1998), social media feeds (Nishtala et al., 2013), and transaction processing services. The workloads are characterized by their burstiness and high concurrency, with millions of small, independent tasks that require rapid response times.
@@
ML fleets combine the scale of WSC workloads with the computational intensity of HPC workloads. They consist of workloads which are focused on running deep neural networks (Krizhevsky et al., 2012), such as recommendation models (Naumov et al., 2019; Zhao et al., 2022) for e-commerce and social media, computer vision tasks for autonomous vehicles (Janai et al., 2020) and medical imaging (Esteva et al., 2021), and natural language processing (NLP) applications like large language models for chatbots and translation services (Zhao et al., 2021).
ML workloads involve heavy matrix-matrix (Fatahalian et al., 2004) or matrix-vector (Bell and Garland, 2008) operations and parallel processing that often require high computational throughput and memory bandwidth. ML workloads also tend to be more communication-bound rather than compute-bound (Li et al., 2014), especially with large models since they are often sharded across many chips and require large datasets to be fed into the model.
ML fleets also face another key challenge, which is rapidly fluctuating user demand and huge tectonic shifts in ML model architectures, as discussed in Section 1. As a result, the designers of an ML fleet need to trade off specialization with fungibility to adapt quickly while still delivering peak performance.
@@ 3. Anatomy of an ML Fleet
In this section, we dive into the anatomy of a production ML fleet to provide perspective on the complexity of managing it. We begin by dissecting the fleet based on its distinctive components and characteristics, starting with the hardware foundation and progressing to the user application level. Figure 3 shows the various layers that comprise the system stack and mediate user access to the ML fleet. This stack illustrates the intricate ecosystem that underpins modern ML operations.
We supplement our discussion in this section with data from a snapshot of Google’s TPU fleet for internal workloads, providing concrete examples of the challenges these systems face in practice. By examining actual usage patterns, resource allocation, and performance metrics from a production ML fleet, we can ground our discussion in real-world scenarios and offer insights based on empirical evidence. This data-driven approach will allow us to illustrate the complexities of managing large-scale ML operations and demonstrate how theoretical concepts translate into practical challenges and opportunities for optimization.
@@ 3.1. Accelerators
ML fleets are distinguished from other types of large-scale compute systems by their accelerator-centric architecture. The ML computing landscape is dominated by domain-specific hardware, such as GPUs and other ASICs. In order to tailor to the vector and matrix intensive operations that underpin ML workloads, new accelerators such as Google’s Tensor Processing Units (TPUs) (Jouppi et al., 2018) have been developed. In general, there has been a Cambrian explosion of ML hardware accelerators (Hennessy and Patterson, 2019), with new accelerators being deployed at an unprecedented rate compared to traditional WSC fleets (Jouppi et al., 2021). Figure 1 vividly illustrates this dynamism, revealing dramatic shifts in our ML fleet’s hardware makeup for internal workloads over just a few years.
@@
Moreover, the heterogeneity extends beyond just accelerator type. Even within a single class of hardware accelerators, there are many different versions of the hardware, adding another layer of complexity to fleet management. Each hardware generation introduces unique features that require significant optimizations to extract peak ML workload efficiency. One notable example is the integration of the SparseCore (SC) in TPUv4 (Jouppi et al., 2023), which was designed to significantly boost performance for embedding-heavy models. Subsequent large-embedding model teams would likely then consider the hardware specifications of the SparseCore when designing their embedding configurations. Design points such as embedding dimension, vocabulary size, valence, and others might also be co-designed to optimize performance on the hardware platform. This demonstrates how hardware-software co-design is becoming increasingly important in improving the efficiency of these diverse accelerators, forming a symbiotic relationship where the computational needs of future workloads affect the next generation of hardware, and the hardware capabilities inform the types of workloads that the ML fleet is best equipped to handle (Shi et al., 2020).
@@ 3.2. Scheduler
The scheduler directly manages the hardware in a fleet by coordinating the allocation of resources; for the case study presented in this paper, it coordinates TPU allocations for Google’s internal-facing ML workloads. There are two interconnected challenges that a scheduler must address when allocating hardware for an ML fleet: (1) optimizing performance across various hardware types, and (2) balancing utilization with stability and fault tolerance.
@@ 3.4. Framework
In addition, ML frameworks often provide auxiliary services to improve efficiency of the entire ML fleet. For instance, TensorFlow’s tf.data (Murray et al., 2021) service optimizes the performance of the data pipeline. These features, while abstracted from the user, can impact the overall system efficiency, as shown in Figure 5. Underneath these high-level frameworks lies a foundation of general-purpose libraries and datacenter services. Frameworks like TensorFlow utilize libraries such as gRPC (Talwar, 2016), protobuf (Varda, 2008), and tcmalloc (Sanjay Ghemawat, [n. d.]) for various low-level operations, and interface with datacenter services for storage (e.g., Colossus (Ghemawat et al., 2003) (Hildebrand and Serenyi, 2021)) and monitoring (e.g., Monarch (Adams et al., 2020)).
@@ 3.5. ML Model & Data
To characterize the ML Fleet at the highest level of the stack, we generally want to know: What types of workloads are we spending most of our compute cycles on? This is a critical question because it drives nearly every design decision we make for the ML Fleet at every level of the stack, from the hardware (how many training vs. inference vs. other chips) to the software (JAX vs. other frameworks, runtime distribution strategies, and compiler optimizations). Workload heterogeneity analysis is useful for understanding what kind of models are prevalent in the production fleet, especially since different workloads stress the hardware in different ways. This understanding can drive decisions about which accelerators to deploy and how many, or which compiler optimizations to carry out.
@@ 4. ML Productivity Goodput
The optimization of ML fleet efficiency is a complex, cyclical challenge, as illustrated in Figure 3. The first challenge is measuring, understanding, and reporting fleetwide efficiency, establishing a baseline for current performance. Second, we must identify and quantify fleet-wide inefficiencies, pinpointing areas that require improvement. Third, we must eliminate these inefficiencies by implementing changes across the fleet, which in turn leads back to the first stage as we measure the impact of these changes. This cycle ensures ongoing optimization and adaptation to the ever-evolving landscape of ML workloads and hardware. To keep pace with this lifecycle, we require a metric that not only quantifies current performance but also guides future optimization efforts across the fleet, which is why Google has developed the ML Productivity Goodput metric.
@@
4.1. Pitfalls & Myths of Traditional Metrics
Before we set the stage for the new metric, we examine the common pitfalls of historical approaches for fleetwide measurement. Figure 9 illustrates how computer architects have historically tended to think about performance metrics (Li et al., 2023; Mars et al., 2011; Kanev et al., 2015). These traditional performance metrics can sometimes fall short in providing a holistic view, given the unique challenges we have discussed in Section 2 and Section 3.
Myth 1: High Capacity equates to high resource availability. While capacity can tell us how many individual accelerators may be available in the fleet at a given time, it does not take into account the topological shape of those accelerators. For example, an ML training workload requesting thousands of chips in a certain physical mesh shape may never be scheduled if the only available accelerators are fragmented across different clusters or data centers. Other factors, such as the geographical location of data storage cells and accelerators, are not included in the capacity metric, even though they significantly affect job scheduling. Therefore, high capacity by itself as a metric does not necessarily effectively translate to high availability for workloads, and we should instead opt for scheduling efficiency as a more robust metric.
Myth 2: High Occupancy guarantees productivity. Occupancy is defined as the fraction of accelerators allocated to jobs and is often measured by the scheduler (e.g. Borg (Verma et al., 2015)). Occupancy is traditionally seen as a key efficiency indicator, but it can be misleading as it masks inefficiencies in the system stack. For example, an accelerator might be successfully allocated but stuck in I/O wait or running poorly optimized code, thus resulting in a high occupancy but very little actual progress being made towards the workload task. This is important for long-running tasks such as ML model training, where frequent pre-emptions may hinder checkpoint progress but still result in a nominally high occupancy. The traditional occupancy metric therefore does not distinguish between productive and unproductive use of allocated resources.
Myth 3: Duty Cycle accurately represents useful work. Duty Cycle measures whether an accelerator is in use, not how much of its compute capacity is used. When looking at an ML workload running on a TPU, duty cycle does not provide any signal on how much the matrix-multiply units (MXUs) are utilized (Jouppi et al., 2023). It is agnostic of the program-level efficiency and does not take into account the effectiveness of the operations being performed. An accelerator could have a high duty cycle while executing unnecessary or redundant computations. So, we require a more sophisticated metric.
Figure 9. Historical utilization-based fleet efficiency metrics. We propose replacing this approach and using goodput as a measure of fleet efficiency rather than utilization.
The Overarching Misconception: Utilization Equals Productivity. The common thread among these metrics is the assumption that keeping accelerators busy equates to productive work. However, this overlooks critical factors. (1) Quality of Computations: None of these metrics assess whether the operations being performed are actually contributing to the desired output. (2) Workload Efficiency: They do not consider whether the workloads are optimally designed for the hardware. (3) System-level Bottlenecks: Focusing solely on accelerator usage ignores potential bottlenecks in data loading, memory access, or inter-accelerator communication. (4) Forward Progress: These metrics provide no insight into how much useful work is being accomplished towards completing an actual ML task.
@@ 4.3. A New Approach: ML Productivity Goodput
This multi-layered structure, as illustrated in Figure 10, offers several advantages over the traditional metrics. First, it allows for precise identification of performance bottlenecks or improvements at specific layers of the stack, facilitates a more granular analysis of efficiency trends over time, and mitigates the risk of misleading interpretations that can arise from aggregated data, such as Simpson’s paradox. [^1] A statistical phenomenon where a trend that is evident within individual groups disappears or reverses when the population groups are combined. Second, by decoupling these submetrics, we enable more targeted optimization efforts and gain deeper insights into the complex interactions within the ML fleet. Finally, this approach not only enhances our ability to measure current performance but also provides a framework for guiding improvements, discussed in Section 5.
@@
Scheduling Goodput (SG) quantifies the efficiency of resource allocation in an ML fleet. It measures the fraction of time that an application has all the required resources simultaneously available to make progress. This metric can be lower than traditional Occupancy, particularly in distributed, bulk-synchronous applications where all required chips must be available concurrently. The numerator of SG is calculated as the simultaneous uptime of all tasks in a distributed ML application that must be connected to make synchronous progress, as shown in Figure 11. This is referred to as “allocated chip-time” or “all-allocated” time. The denominator is fleet capacity, expressed as chip-time. This provides a full view of how effectively the scheduling layer is using the fleet’s resources.

### v2

@@ Abstract.
Machine learning (ML) infrastructures operating at warehouse scale present unique performance characterization challenges beyond traditional high-performance computing metrics. This paper introduces a systematic framework for analyzing ML fleet efficiency, demonstrated on Google’s production TPU infrastructure comprising thousands of accelerators running diverse workloads. Our fleet-wide analysis reveals performance dependencies spanning the entire ML system stack, from hardware to model architecture, data pipelines, frameworks, compilers, and schedulers. We identify critical gaps in conventional utilization-based performance metrics and propose ”ML Productivity Goodput” (MPG) to capture fleet-wide efficiency across heterogeneous ML environments. Using MPG, we provide different methods for identifying and addressing bottlenecks throughout the ML system stack. When applied to Google’s TPU workloads that are used in production, our results show that the methodology yields substantial performance improvements, and establishes a basis for efficient management of large-scale ML computing infrastructure.
@@ 1. Introduction
This paper presents a playbook for instrumenting and optimizing large-scale ML fleets, using Google’s production Tensor Processing Unit (TPU) infrastructure for internal workloads as our case study. In contrast to prior work, where previous Google TPU papers have predominantly focused on hardware architecture design and chip-level performance evaluations (Jouppi et al., 2023, 2021, 2017), our work provides the first in-depth perspective on the software stack that enables these domain-specific accelerators to operate efficiently at scale.
[…]
2 Our second contribution addresses these issues. To view the fleet holistically and capture its performance across multiple layers of the ML system stack, we introduce the “ML Productivity Goodput” (MPG) metric. Unlike traditional metrics, MPG encompasses scheduling efficiency, runtime efficiency, and compiler / program efficiency. MPG is analogous to the “iron law” of computer performance (Eeckhout, 2010), but adapted for ML fleets as shown in Figure 2. The MPG metric enables us to identify areas for improvement across the entire machine learning fleet by analyzing performance across different segments, such as accelerator type, model architecture, and workload phase (e.g., training vs. serving). By breaking down the metric into subcomponents—Scheduling Goodput (SG), Runtime Goodput (RG), and Program Goodput (PG)—and examining it along the axes of fleet characteristics, we can precisely pinpoint optimization opportunities.
[…]
3 Finally, we provide a procedure for leveraging the MPG metric and show strategies that we deployed in the real world for optimizing ML fleet performance with the help of this metric. We show how MPG helps us identify compiler optimizations to generate more efficient code for specific hardware configurations, and improve scheduling algorithms to better utilize heterogeneous resources within the fleet, along with a few other strategies. Even after optimizations are implemented in the fleet, the MPG metric can be used to validate and track fleet-wide improvements.
Figure 2. ML Productivity Goodput (MPG) and its components.
@@ 2.1. Workloads
WSCs (Barroso and Hölzle, 2009) traditionally focus on Internet-scale services, which serve web content, process user queries, and handle data storage and retrieval for billions of users. WSCs must manage diverse, often latency-sensitive tasks such as search indexing (Brin and Page, 1998), social media feeds (Nishtala et al., 2013), and transaction processing services. The workloads are characterized by their burstiness and high concurrency, with millions of small, independent tasks that require fast response times.
@@
ML fleets combine the scale of WSC workloads with the computational intensity of HPC workloads. They consist of workloads which are focused on running deep neural networks (Krizhevsky et al., 2012), such as recommendation models (Naumov et al., 2019; Zhao et al., 2022) for e-commerce and social media, computer vision tasks for autonomous vehicles (Janai et al., 2020) and medical imaging (Esteva et al., 2021), and natural language processing (NLP) applications like large language models for chatbots and translation services (Zhao et al., 2021). ML workloads involve heavy matrix-matrix (Fatahalian et al., 2004) or matrix-vector (Bell and Garland, 2008) operations and parallel processing that often require high computational throughput and memory bandwidth. ML workloads also tend to be more communication-bound rather than compute-bound (Li et al., 2014), especially since large models are often sharded across many chips and require large datasets to be fed as input. ML fleets also face another key challenge, which is rapidly fluctuating user demand and huge tectonic shifts in ML model architectures, as discussed in Section 1. As a result, the designers of an ML fleet need to trade off specialization with fungibility to adapt quickly while still delivering peak performance.
@@ 3. Anatomy of an ML Fleet
We begin by dissecting Google’s user-facing production ML fleet, analyzing its components from hardware foundation to application level. Figure 3 shows the layers comprising the system stack that mediate user access. We supplement our discussion with data from a snapshot of Google’s TPU fleet for internal workloads, providing concrete examples of challenges these systems face.
@@ 3.1. Accelerators
ML fleets are distinguished from other types of large-scale systems by their accelerator-centric architecture. The ML computing landscape is dominated by domain-specific hardware, such as GPUs and other ASICs. In order to tailor to the vector and matrix intensive operations that underpin ML workloads, there has been a Cambrian explosion of ML hardware accelerators (Hennessy and Patterson, 2019), with new accelerators being deployed at an unprecedented rate compared to traditional WSC fleets (Jouppi et al., 2021). Figure 1 illustrates this dynamism, showing dramatic shifts in our ML fleet’s hardware makeup for internal workloads over just a few years.
@@
Moreover, the heterogeneity extends beyond just accelerator type. Even within a single class of hardware accelerators, there are many different versions of the hardware, adding another layer of complexity to fleet management. Each hardware generation introduces unique features that require significant optimizations to extract peak ML workload efficiency. One notable example is the integration of the SparseCore (SC) in TPUv4 (Jouppi et al., 2023), which was designed to significantly boost performance for embedding-heavy models. Subsequent large-embedding model teams would likely then consider the hardware specifications of the SparseCore when designing their embedding configurations. Model design configurations such as embedding dimension, vocabulary size, valence, and others might also be co-designed to optimize performance on the hardware platform. This shows how hardware-software co-design is becoming increasingly important in improving the efficiency of these diverse accelerators, forming a symbiotic relationship where the computational needs of future workloads affect the next generation of hardware, and the hardware capabilities inform the types of workloads that the ML fleet is best equipped to handle (Shi et al., 2020).
@@ 3.2. Scheduler
The scheduler directly manages the hardware in a fleet by coordinating the allocation of resources. For the case study presented in this paper, it coordinates TPU allocations for Google’s internal-facing ML workloads. There are two interconnected challenges that a scheduler must address when allocating hardware for an ML fleet: (1) optimizing performance across various hardware types, and (2) balancing utilization with stability and fault tolerance.
@@ 3.4. Framework
Also, ML frameworks often provide auxiliary services to improve efficiency of the entire ML fleet. For instance, TensorFlow’s tf.data (Murray et al., 2021) service optimizes the performance of the data pipeline. These features, while abstracted from the user, can impact the overall system efficiency, as shown in Figure 5. Underneath these high-level frameworks lies a foundation of general-purpose libraries and datacenter services. Frameworks like TensorFlow utilize libraries such as gRPC (Talwar, 2016), protobuf (Varda, 2008), and tcmalloc (Sanjay Ghemawat, [n. d.]) for various low-level operations, and interface with datacenter services for storage (e.g., Colossus (Ghemawat et al., 2003) (Hildebrand and Serenyi, 2021)) and monitoring (e.g., Monarch (Adams et al., 2020)).
@@ 3.5. ML Model & Data
To characterize the ML Fleet at the highest level of the stack, we want to know: What types of workloads are we spending most of our compute cycles on? This is a critical question because it drives nearly every design decision we make for the ML Fleet at every level of the stack, from the hardware (how many training vs. inference vs. other chips) to the software (JAX vs. other frameworks, runtime distribution strategies, and compiler optimizations). Workload heterogeneity analysis is useful for understanding what kind of models are prevalent in the production fleet, especially since different workloads stress the hardware in different ways. This understanding can drive decisions about which accelerators to deploy and how many, or which compiler optimizations to carry out.
@@ 4. ML Productivity Goodput
The optimization of ML fleet efficiency is a complex, cyclical challenge, as illustrated in Figure 3. The first challenge is measuring, understanding, and reporting fleetwide efficiency, establishing a baseline for current performance. Second, we must identify and quantify fleet-wide inefficiencies, pinpointing areas that require improvement. Third, we must eliminate these inefficiencies by implementing changes across the fleet, which in turn leads back to the first stage as we measure the impact of these changes. This cycle ensures ongoing optimization and adaptation to the ever-evolving landscape of ML workloads and hardware. To keep pace with this lifecycle, we require a metric that not only quantifies current performance but also guides optimization efforts across the fleet.
@@ 4.2. A New Approach: ML Productivity Goodput
This multi-layered structure, as shown in Figure 9, offers several advantages over traditional metrics. First, it allows precise identification of performance bottlenecks or improvements at specific layers of the stack, facilitates a more granular analysis of efficiency trends over time, and mitigates the risk of misleading interpretations that can arise from aggregated data. Second, by decoupling these submetrics, we enable more targeted optimization efforts and gain deeper insights into the complex interactions within the ML fleet. Finally, this approach not only enhances our ability to measure current performance but also provides a framework for guiding improvements, discussed in Section 5.
@@
Scheduling Goodput (SG) quantifies the efficiency of resource allocation in an ML fleet. It measures the fraction of time that an application has all the required resources simultaneously available to make progress. This metric can be lower than traditional Occupancy metrics, particularly in distributed, bulk-synchronous applications where all required chips must be available concurrently. The numerator of SG is calculated as the simultaneous uptime of all tasks in a distributed ML application that must be connected to make synchronous progress, as shown in Figure 10. This is referred to as “allocated chip-time” or “all-allocated” time. The denominator is fleet capacity, expressed as chip-time. This provides a full view of how effectively the scheduling layer is using the fleet’s resources.